# TPV: Parameter Perturbations Through the Lens of Test Prediction Variance

### v2

@@ Abstract
Theoretically, we show that TPV estimated on the training set converges to its test-set value in the overparameterized limit, providing the first result that prediction variance under local parameter perturbations can be inferred from training inputs alone, and this stability is decoupled from generalization performance. Empirically, TPV exhibits a striking stability across datasets and architectures—including extremely narrow networks— and TPV correlates well with clean test loss.
Finally, we demonstrate that modeling pruning as a TPV perturbation yields a simple label-free importance measure that performs competitively with state-of-the-art pruning methods, illustrating the practical utility of TPV. Code Here
@@ 1 Introduction
Figure 1: TPV stability on synthetic data: Each point corresponds to one synthetic configuration (dataset type, input dimension, network width, depth) and one perturbation source (label noise; SGD noise). Axes show empirical TPV on the training and test sets; $y=x$ is the ideal reference line along with a gray colored $50\%$ error band; colormap indicates generalization gap $L_{\text{test}}-L_{\text{train}}$. We ran 324 configurations from label-noise and SGD-noise, spanning more than five orders of magnitude in TPV and different levels of generalization gaps. Despite this heterogeneity, all points concentrate tightly around the diagonal $\mathrm{TPV}_{\text{train}}=\mathrm{TPV}_{\text{test}}$; most surprisingly, even for width=1=1. This demonstrates that: i) TPV stability holds true even at extremely low widths; ii) TPV stability is decoupled from generalization. See section 5.1 for more experiments.
@@
    Correlation with Test Loss: We find that empirical TPV estimates correlate well with test loss. Together with TPV stability, it provides a new way to gauge generalization and robustness using training dataset at a granular level by plugging in specific noise mechanisms.
@@
    Applications to label noise, SGD noise, and pruning. We derive TPV expressions for label noise and SGD stationary noise, reveal how overparameterization and wide minima suppresses prediction variance, and introduce JBR—a practical pruning criterion derived from TPV geometry—that matches or exceeds state-of-the-art baselines on CIFAR-10/100 and ImageNet.
@@ 2 Test Prediction Variance
This local perspective is more directly appropriate for understanding robustness to realistic noise sources near or after convergence—label noise at convergence, SGD stationary noise, quantization noise, dropout during inference, and post-training pruning masks—all of which act as perturbations around a trained model rather than as full retraining procedures.
Crucially, TPV conditions on a particular trained solution $w^{\star}$, whereas global variance averages over the entire distribution of solutions obtainable by the learning algorithm. The two quantities therefore capture fundamentally different notions of variability.
@@ 2.1 Our starting point: expected test error variance
Let $\delta w$ denote a zero-mean parameter perturbation drawn from a distribution $R$ that models the sources of noise acting near or after convergence (label noise at convergence, SGD stationary noise, finite precision, pruning masks, etc.), denote its covariance matrix by $C:=\mathbb{E}_{R}[\delta w\,\delta w^{\top}]\in\mathbb{R}^{p\times p}$. The perturbed predictor is $f_{w^{\star}+\delta w}$. For a test pair $(x,y)$ with $y=f^{\star}(x)$, the expected test error is
@@
is the Jacobian of the model output with respect to parameters at $w^{*}$.
Under this approximation,
@@
Eq. (6) is the central object in this paper: TPV, and in turn the expected test error (due to Eq. (3)), is controlled by the interaction between the curvature spectrum of $H_{\mathrm{eff}}$ and the perturbation covariance $C$.
@@ 2.2 TPV Trace Stability
In this section, we prove that for trained overparameterized networks of width $m\rightarrow\infty$, TPV estimated entirely using the training dataset acts as a good estimator of the test set TPV object, irrespective of a model’s generalization performance.
@@
where $c_{1}:=(2\varepsilon_{\mathrm{NTK}}+o_{n_{\mathrm{tr}},n_{\mathrm{te}}}(1))$ and $C$ is the parameter perturbation covariance matrix, such that $\varepsilon_{\mathrm{NTK}}\rightarrow 0$ as $m\rightarrow\infty$ and $o_{n,n_{\mathrm{te}}}(1)\rightarrow 0$ as $n,n_{\mathrm{te}}\to\infty$.
@@
The intuition behind the proof is simple and uses two observations–1. Jacot et al. (2018); Allen-Zhu et al. (2019) show that the NTK kernel remains stable during training; 2. due to random initialization and i.i.d. sampling of datasets, $H_{\mathrm{eff}}(w_{0};X)$ concentrates around its population counterpart in operator norm for both $X_{\mathrm{tr}}$ and $X_{\mathrm{te}}$ due to the Law of Large Numbers. Combining these two observations allows us to upper-bound the distance between the training set and test set TPV objects. We use this assumption in section 3.2 and 3.3.
@@ 3 Parameter Perturbation Sources
In this section, we study the effect of individual noise sources in isolation on test prediction variance (TPV). Specifically, we revisit label noise, SGD mini-batch noise, and finite-precision noise, and pruning noise and show how the benefits of over-parameterization and wide minima arise from the common mechanism of suppressing TPV. We also model pruning as a TPV perturbation and propose an algorithm that performs competitively with the state-of-the-art baselines.
@@ 3.1 Label Noise
In this section, we do not invoke the TPV trace stability result and study the TPV object in its general form.
@@
The proof is shown in Appendix B for completeness. Note that in the linear regression setting above, the “global” prediction variance $\mathbb{E}_{x}\operatorname{Var}_{\varepsilon}(f_{w^{\star}}(x))$ obtained by retraining on different noisy label realizations coincides exactly with the local TPV quantity $\mathrm{Tr}(H_{\mathrm{eff}}C)$, and is a well known result (Bartlett et al., 2020) in literature characterizing benign overfitting in linear regression models. This is because the conditioning of the matrix $XX^{\top}$ improves as $d$ grows beyond $n$, resulting in lower $\operatorname{TPV}_{\mathrm{label}}$, making the model less sensitive to label noise. Even more specifically, if training and test distributions are identical, then for a large enough $n(\ll d)$, we have $\frac{1}{n}\,X^{\top}X~=I_{d}$ (due to the whitening assumption). Under isotropic/whitened assumptions on the rows of $X$, random matrix theory (Wishart) gives $\mathbb{E}(XX^{\top})^{-1}\approx\frac{1}{d}I_{n}$. Therefore, $\operatorname{TPV}_{\mathrm{label}}\approx\sigma_{\varepsilon}^{2}n/d$.
@@
The proof is provided in Appendix C. Equation 10 expresses the contribution of label noise to test prediction variance as a spectral sum involving (i) test-distribution Jacobian modes $B_{ii}$ and (ii) the corresponding training singular values $s_{i}$. The linear result (Eq. 9) is a special case of Eq. 10 when the nonlinear Jacobian is replaced by the data matrix. To our knowledge, Eq. 10 is the first expression of label-noise–induced prediction variance for finite-width deep networks in terms of a test-Jacobian operator and a general parameter-noise covariance $C$, that also accommodates other perturbation sources (e.g., SGD noise, finite precision, pruning masks) within the same $\operatorname{Tr}(H_{\mathrm{eff}}C)$ template.
Equation 10 highlights that label-noise sensitivity is dominated by directions where $B_{ii}$ is large while $s_{i}$ is small—i.e., where the test-distribution Jacobian aligns with poorly conditioned training directions. This makes explicit how the local geometry around $w^{\star}$—particularly the conditioning and alignment of Jacobian modes—governs test-prediction variance.
@@
Two pathological regimes are worth noting w.r.t. Eq. 10. (i) When the linearized system around $w^{\star}$ interpolates the noisy labels (e.g., $J\delta w=\varepsilon$ with full row rank), both the training TPV and the label-noise TPV in Eq. 10 collapse to $\sigma_{\varepsilon}^{2}$ (or to 0 as $\sigma_{\varepsilon}^{2}\!\to\!0$) if the training and test distributions match. TPV stability then holds trivially, but TPV becomes uninformative across architectures. (ii) When $\sigma_{\varepsilon}^{2}$ is sufficiently large that the noisy solution either leaves the linearization regime (breaking the assumptions of Theorem 3.2) or induces a large $\mathrm{Tr}(C)$ making the upper bound in Theorem 2.1 weak, it can both cause TPV stability to break. The non-degenerate, small-but-finite noise regime—where interpolation does not occur and sufficiently small perturbations are used, and the min-norm solution is achieved—is the regime in which label noise TPV exhibits meaningful geometric dependence and is useful. On a related note, SGD in deep networks may not find the min-norm solution in $\delta w$ and is a proxy for the true optimization problem. See Appendix C.2 and C.3 for a detailed discussion on the label noise TPV object.
@@ 3.2 SGD Stationary Noise Near Convergence
Consider a scalar-output non-linear model $f(x;w)$ trained using SGD with squared loss $L$ on a fixed dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ sampled i.i.d. from an underlying distribution $\mathcal{D}$ without any label noise, and assume that there exists a minimizer $w^{\star}$ that achieves zero training loss. Then,
@@
The proof is shown in Appendix D and invokes the TPV trace stability result because we use $H_{\mathrm{eff}}$ computed on the training set.
@@ 3.3 Finite-precision noise
Consider a scalar-output non-linear model $f(x;w)$ trained using GD (not SGD) with squared loss $L$ on a fixed dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ sampled i.i.d. from an underlying distribution $\mathcal{D}$ without any label noise, and assume that there exists a minimizer $w^{\star}$ that achieves zero training loss. Then,
@@
The proof is shown in Appendix E. Thus, similar to section 3.2, the above claim shows that sharper minima increase generalization error under quantization/finite precision.
3.4 Jacobian-Based Rebalancing (JBR) for Pruning
$u(x;w):=-\log(p_{c(x)}(x;w)),$
Pruning as a Source of Parameter Noise
$\delta w_{g}=\sigma\,\xi_{g}\,w_{g},\qquad g\in\mathcal{G},$ (13)
Connection and Contrast with JC
@@ 4.1 Test Prediction Variance
Work on the Neural Tangent Kernel (NTK) shows that, in the infinite-width limit, gradient descent training of neural networks is equivalent to kernel regression with a fixed kernel, yielding precise characterizations of training dynamics and generalization in terms of the NTK spectrum (Jacot et al., 2018; Lee et al., 2019; Arora et al., 2019). Our TPV framework leverages this infinite-width NTK regime to obtain a trace-stability theorem —showing that $\mathrm{Tr}\bigl(H_{\mathrm{eff}}\,C_{\mathrm{prune}}\bigr)$ computed on training set converges to the test set estimate—while providing a curvature–covariance interpretation of prediction variance applicable to finite networks as well.
4.2 Pruning
@@ 5 Experiments
We run experiments around TPV stability between training and test sets and empirically identify the conditions when it holds and when it breaks. We then study TPV under label noise and how width plays an important role in modulating it. We do not include experiments for SGD based perturbations and the relation between the flatness of minima and TPV because several prior research works have confirmed a similar relationship (Keskar and et al., 2017; Smith and Le, 2018; Jastrzebski et al., 2017). Finally, we run experiments on the proposed JBR pruning algorithm comparing its performance to existing post-training pruning methods.
@@ 5.1 TPV Stability
To stress-test TPV trace stability beyond any single model or dataset, we construct a large synthetic benchmark spanning a broad grid of architectures, data distributions, and perturbation sources. We perform two sets of experiments– in the first experiment, we fix network width in $\{1,256\}$ and for each width experiment with all combinations of the other configurations; in the second experiment, we fix the number of training samples in $\{10,10000\}$ and do the same. In each set, there are a total of $324$ distinct experiments (each with $20$ independent runs). For each combination, we first train a clean reference network $f_{w^{\star}}$ on noiseless regression targets. Around each such $f_{w^{\star}}$, we then instantiate two different parameter-noise mechanisms– label noise and SGD stationary noise– and train independent copies of the reference network under these noise mechanisms and get a perturbed version of the network. Importantly, we discard the runs that violate the first order assumption (using a simple finite-difference approximation check) and in which noisy model loss does not go down. For full details, see Appendix F.1.
For each configuration we estimate the empirical TPV on the training set $X_{\text{tr}}$ and test set $X_{\text{te}}$,
@@
We now verify TPV stability on CIFAR-10 dataset. Similar to the synthetic data experiment above, we first use architectures of different widths and fix the number of training samples. The scatter plot is shown in Fig. 3. Similar to Fig. 1, TPV remains stable across different widths.
We next study the effect of the number of training samples. The scatter plot is shown in Fig. 4. Once again, we find that TPV stability breaks when $n_{\text{train}}$ is extremely low ($n_{\text{train}}=1$ lies outside the $50\%$ error band), holds better for $n_{\text{train}}=10$ (mostly within the band), and very tightly for $n_{\text{train}}=10000$. Analogous experiments on CIFAR-100 are shown in Fig. 12 and 12 in the Appendix. Full experimental details are in Appendix F.1.2.
Interpretation
: We make three important observations from these experiments: i) points corresponding to large generalization gaps are just as close to the diagonal as those with small gaps as long as we use a large $n_{\text{train}}$. This shows that empirical TPV computed using training and test sets match closely irrespective of the generalization performance of the model. We call this TPV stability [^1] TPV stability is different from TPV trace stability since former uses Eq. 2 while the latter uses Eq. 6.. Therefore, TPV stability is decoupled from generalization; ii) TPV stability holds even when network width is very small; iii) TPV stability breaks when the number of samples used for estimation is too small.
@@
Figure 7: Empirical and theoretical TPV estimates under label noise on synthetic data for noise standard deviation $\sigma=0.1$. As width increases, theoretical TPV and empirical test set TPV reduce, but training set TPV increases, breaking TPV stability when $\sigma$ is large.

### v3

@@ Abstract
Theoretically, we show that TPV estimated on the training set converges to its test-set value in the overparameterized limit, providing the first result that prediction variance under local parameter perturbations can be inferred from training inputs alone, and this stability is decoupled from generalization performance. Empirically, TPV exhibits a striking stability across datasets and architectures even for extremely narrow networks. Further, TPV correlates well with test loss, serving as a training-set based predictive metric for generalization. Code Available Here
@@ 1 Introduction
Figure 1: TPV stability on synthetic data: Each point corresponds to one synthetic configuration (dataset type, input dimension, network width, depth) and one perturbation source (label noise; SGD noise). Axes show empirical TPV on the training and test sets; $y=x$ is the ideal reference line along with a gray colored $50\%$ error band; colormap indicates generalization gap $L_{\text{test}}-L_{\text{train}}$. We ran 324 configurations from label-noise and SGD-noise, spanning more than five orders of magnitude in TPV and different levels of generalization gaps. Despite this heterogeneity, all points concentrate tightly around the diagonal $\mathrm{TPV}_{\text{train}}=\mathrm{TPV}_{\text{test}}$; most surprisingly, even for width=1=1. This demonstrates that: i) TPV stability holds true even at extremely low widths; ii) TPV stability is decoupled from generalization.
@@
    Correlation with Test Loss: We find that empirical TPV estimates correlate well with test loss. Together with TPV stability, it provides a new post-training diagnostic tool using training dataset for characterizing model robustness and generalization behavior under realistic perturbations without access to test labels, making TPV a practical diagnostic for deployed models.
@@
    Applications to label noise, SGD noise, and pruning: We derive TPV expressions for label noise and SGD stationary noise, reveal how overparameterization and wide minima suppresses prediction variance, and introduce JBR—a practical pruning criterion derived from TPV geometry—that matches or exceeds state-of-the-art baselines on CIFAR-10/100 and ImageNet.
@@ 2 Test Prediction Variance
This local perspective is more directly appropriate for understanding robustness to realistic noise sources near or after convergence—label noise at convergence, SGD stationary noise, quantization noise, dropout during inference, and post-training pruning masks—all of which act as perturbations around a trained model rather than as full retraining procedures. In contrast, global variance averages over the entire distribution of solutions obtainable by the learning algorithm. The two quantities therefore capture fundamentally different notions of variability.
@@ 2.1 Our starting point: expected test error variance
Let $\delta w$ denote a zero-mean parameter perturbation drawn from a distribution $R$ that models the sources of noise acting near or after convergence (label noise at convergence, SGD stationary noise, finite precision, pruning masks, etc.), denote its covariance matrix by $C:=\mathbb{E}_{R}[\delta w\,\delta w^{\top}]\in\mathbb{R}^{p\times p}$. The perturbed predictor is $f_{w^{\star}+\delta w}$. For a test pair $(x,y)$ with $y=f^{\star}(x)$ (noiseless labels), the expected test error is
@@
is the Jacobian of the model output with respect to parameters at $w^{*}$. Under this approximation,
@@
Eq. (6) is the central object in this paper: TPV, and in turn the expected test error (due to Eq. (3)), is controlled by the interaction between the curvature spectrum of $H_{\mathrm{eff}}$ and the perturbation covariance $C$. See Appendix B for TPV equation under scalar vs vector output models.
@@ 2.2 TPV Trace Stability
In this section, we prove that for trained overparameterized networks of width $m\rightarrow\infty$, TPV estimated entirely using the training dataset acts as a good estimator of the test set TPV object, irrespective of a model’s generalization performance. The analysis relies on simplifying assumptions (e.g., isotropic parameter perturbations) that enable theoretical tractability and serve to motivate the empirical TPV stability observed in more general settings.
@@
where $c_{1}:=\frac{(n_{\mathrm{tr}}+n_{\mathrm{te}})}{p}\,\varepsilon_{\mathrm{NTK}}~+~o_{n_{\mathrm{tr}},n_{\mathrm{te}}}(1)$ and $C$ is the parameter perturbation covariance matrix (assumed to be isotropic), such that $\varepsilon_{\mathrm{NTK}}\rightarrow 0$ as $m\rightarrow\infty$ and $o_{n,n_{\mathrm{te}}}(1)\rightarrow 0$ as $n,n_{\mathrm{te}}\to\infty$.
@@
The proof uses two observations–1. Jacot et al. (2018); Allen-Zhu et al. (2019) show that the NTK kernel remains stable during training; 2. due to random initialization and i.i.d. sampling of datasets, $H_{\mathrm{eff}}(w_{0};X)$ concentrates around its population counterpart in operator norm for both $X_{\mathrm{tr}}$ and $X_{\mathrm{te}}$ due to the Law of Large Numbers. Combining these two observations allows us to upper-bound the distance between the training set and test set TPV objects.
@@ 3 Parameter Perturbation Sources
In this section, we study the effect of individual noise sources in isolation on test prediction variance (TPV). Specifically, we revisit label noise, SGD mini-batch noise, and finite-precision noise, and show how the benefits of over-parameterization and wide minima arise from the common mechanism of suppressing TPV. We also model pruning as a TPV perturbation and propose an algorithm that performs competitively with the state-of-the-art baselines across CIFAR-10/100 and ImageNet (see Appendix H).
@@ 3.1 Label Noise
In this section, we do not assume TPV stability and study the TPV object in its general form.
@@
The proof is shown in Appendix C for completeness. Note that in the linear regression setting above, the “global” prediction variance $\mathbb{E}_{x}\operatorname{Var}_{\varepsilon}(f_{w^{\star}}(x))$ obtained by retraining on different noisy label realizations coincides exactly with the local TPV quantity $\mathrm{Tr}(H_{\mathrm{eff}}C)$, and is a well known result (Bartlett et al., 2020) in literature characterizing benign overfitting in linear regression models. This is because the conditioning of the matrix $XX^{\top}$ improves as $d$ grows beyond $n$, resulting in lower $\operatorname{TPV}_{\mathrm{label}}$, making the model less sensitive to label noise. Even more specifically, if training and test distributions are identical, then for a large enough $n(\ll d)$, we have $\frac{1}{n}\,X^{\top}X~=I_{d}$ (due to the whitening assumption). Under isotropic/whitened assumptions on the rows of $X$, random matrix theory (Wishart) gives $\mathbb{E}(XX^{\top})^{-1}\approx\frac{1}{d}I_{n}$. Therefore, $\operatorname{TPV}_{\mathrm{label}}\approx\sigma_{\varepsilon}^{2}n/d$.
@@
The proof is provided in Appendix D. Equation 10 expresses the contribution of label noise to test prediction variance as a spectral sum involving (i) test-distribution Jacobian modes $B_{ii}$ and (ii) the corresponding training singular values $s_{i}$. The linear result (Eq. 9) is a special case of Eq. 10 when the nonlinear Jacobian is replaced by the data matrix. To our knowledge, Eq. 10 is the first expression of label-noise–induced prediction variance for finite-width deep networks in terms of a test-Jacobian operator and a general parameter-noise covariance $C$, that also accommodates other perturbation sources (e.g., SGD noise, finite precision, pruning masks) within the same $\operatorname{Tr}(H_{\mathrm{eff}}C)$ template.
Equation 10 highlights that label-noise sensitivity is dominated by directions where $B_{ii}$ is large while $s_{i}$ is small—i.e., where the test-distribution Jacobian aligns with poorly conditioned training directions. This makes explicit how the local geometry around $w^{\star}$—particularly the conditioning and alignment of Jacobian modes—governs TPV.
@@
Two pathological regimes are worth noting w.r.t. Eq. 10. (i) When the linearized system around $w^{\star}$ interpolates the noisy labels (e.g., $J\delta w=\varepsilon$ with full row rank), both the training TPV and the label-noise TPV in Eq. 10 collapse to $\sigma_{\varepsilon}^{2}$ (or to 0 as $\sigma_{\varepsilon}^{2}\!\to\!0$) if the training and test distributions match. TPV stability then holds trivially, but TPV becomes uninformative across architectures. (ii) When $\sigma_{\varepsilon}^{2}$ is sufficiently large that the noisy solution either leaves the linearization regime (breaking the assumptions of Theorem 3.2) or induces a large $\mathrm{Tr}(C)$ making the upper bound in Theorem 2.1 weak, it can both cause TPV stability to break. The non-degenerate, small-but-finite noise regime—where interpolation does not occur and sufficiently small perturbations are used, and the min-norm solution is achieved—is the regime in which label noise TPV exhibits meaningful geometric dependence and is useful. On a related note, SGD in deep networks may not find the min-norm solution in $\delta w$ and is a proxy for the true optimization problem. See Appendix D.2 and D.3 for a detailed discussion on the label noise TPV object.
@@ 3.2 SGD Stationary Noise Near Convergence
Consider a scalar-output non-linear model $f(x;w)$ trained using SGD with squared loss $L$ on a fixed dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ sampled i.i.d. from an underlying distribution $\mathcal{D}$ without any label noise, and assume that there exists a minimizer $w^{\star}$ with small but nonzero residuals. Then,
@@
The proof is shown in Appendix E and assumes TPV stability because we use $H_{\mathrm{eff}}$ computed on the training set.
@@ 3.3 Finite-precision noise
Consider a scalar-output non-linear model $f(x;w)$ trained using GD (not SGD) with squared loss $L$ on a fixed dataset $\{(x_{i},y_{i})\}_{i=1}^{n}$ sampled i.i.d. from an underlying distribution $\mathcal{D}$ without any label noise, and assume that there exists a minimizer $w^{\star}$ with small but nonzero residuals. Then,
@@
The proof is shown in Appendix F. Thus, similar to section 3.2, the above claim shows that sharper minima increase generalization error under quantization/finite precision.
@@ 4 Relation to prior work
Work on the Neural Tangent Kernel (NTK) shows that, in the infinite-width limit, gradient descent training of neural networks is equivalent to kernel regression with a fixed kernel, yielding precise characterizations of training dynamics and generalization in terms of the NTK spectrum (Jacot et al., 2018; Lee et al., 2019; Arora et al., 2019). Our TPV framework leverages this infinite-width NTK regime to obtain a trace-stability theorem —showing that $\mathrm{Tr}\bigl(H_{\mathrm{eff}}\,C\bigr)$ computed on training set converges to the test set estimate— applicable to finite networks.
@@ 5 Experiments
We run experiments around TPV stability between training and test sets and empirically identify the conditions when it holds and when it breaks. We then study TPV under label noise and how width plays an important role in modulating it. We do not include experiments for SGD based perturbations and the relation between the flatness of minima and TPV because several prior research works have confirmed a similar relationship (Keskar and et al., 2017; Smith and Le, 2018; Jastrzebski et al., 2017). Pruning experiments are shown in Appendix H.
@@ 5.1 TPV Stability
To stress-test TPV trace stability beyond any single model or dataset, we construct a large synthetic benchmark spanning a broad grid of architectures, data distributions, and perturbation sources. We perform two sets of experiments– in the first experiment, we fix network width in $\{1,256\}$ and for each width experiment with all combinations of the other configurations; in the second experiment, we fix the number of training samples in $\{10,10000\}$ and do the same. In each set, there are a total of $324$ distinct experiments (each with $20$ independent runs). For each combination, we first train a clean reference network $f_{w^{\star}}$ on noiseless regression targets. Around each such $f_{w^{\star}}$, we then instantiate two different parameter-noise mechanisms– label noise and SGD stationary noise– and train independent copies of the reference network under these noise mechanisms and get a perturbed network. Importantly, we discard the runs that violate the first order assumption (using a finite-difference approximation check) and in which noisy model loss does not go down. For full details, see Appendix G.1. For each configuration we estimate the empirical TPV on the training set $X_{\text{tr}}$ and test set $X_{\text{te}}$,
@@
We now verify TPV stability on CIFAR-10 dataset. Similar to the synthetic data experiment above, we first use architectures of different widths and fix the number of training samples. The scatter plot is shown in Fig. 3. Similar to Fig. 1, TPV remains stable with width.
We next study the effect of the number of training samples. The scatter plot is shown in Fig. 8. Once again, we find that TPV stability breaks when $n_{\text{train}}$ is extremely low ($n_{\text{train}}=1$ lies outside the $50\%$ error band), holds better for $n_{\text{train}}=10$ (mostly within the band), and very tightly for $n_{\text{train}}=10000$. Analogous experiments on CIFAR-100 are shown in Fig. 10 and 10 in the Appendix. Full experimental details are in Appendix G.1.2.
Interpretation:
We make three important observations from these experiments: i) points corresponding to large generalization gaps are just as close to the diagonal as those with small gaps as long as we use a large $n_{\text{train}}$. This shows that empirical TPV computed using training and test sets match closely irrespective of the generalization performance of the model. We call this TPV stability [^1] TPV stability is different from TPV trace stability since former uses Eq. 2 while the latter uses Eq. 6.. Therefore, TPV stability is decoupled from generalization; ii) TPV stability holds even when network width is very small; iii) TPV stability breaks when the number of samples used for estimation is too small.
@@
Figure 6: Empirical and theoretical TPV estimates under label noise on synthetic data for noise $\sigma=0.1$. As width increases, theoretical TPV and empirical test set TPV reduce, but training set TPV increases, breaking TPV stability when $\sigma$ is large.